Through careful record-keeping, you have the times to burn out of 250 Bright brand light bulbs. The average time to burn out is 400 hours, with a standard deviation of 14 hours. Give a 95% confidence interval for the meantime to burn out of this brand of the light bulb. Round your answers to 4 decimal places

1
Expert's answer
2021-12-06T16:09:58-0500

The critical value for "\alpha = 0.05" and "df = n-1 = 249" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} = 1.969537."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}},\bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

"=(400-1.969537\times\dfrac{14}{\sqrt{250}},400+1.969537\times\dfrac{14}{\sqrt{250}})"

"=(398.256,401.744)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "398.256 < \mu < 401.744," which indicates that we are 95% confident that the true population mean "\mu" is contained by the interval "(398.256, 401.744)."